# MInCo: Mitigating Information Conflicts in Distracted Visual Model-based Reinforcement Learning

### v2

@@ Abstract
Continues control under visual distractions, model-based reinforcement learning, information conflicts.
@@ I Introduction
Among various forms of machine learning, reinforcement learning (RL), which models decision-making through trial and error, is one of the approaches most similar to human learning [1]. Deep reinforcement learning (DRL) utilizes deep neural networks to tackle high-dimensional tasks, achieving notable success in fields such as robotic manipulation [3, 4], autonomous driving [5], video games [6, 7], Go [8], and others [9]. Another fundamental human ability is planning actions based on predictions of the environment, which is especially useful in long-term tasks. Model-based reinforcement learning (MBRL), designed to mimic this ability, learns a dynamics model to predict the future and then performs policy optimization [10, 11, 12] or planning [13, 14, 15, 16] to determine the optimal action sequence. However, modeling environmental dynamics from high-dimensional visual observations remains challenging, particularly when faced with noise and distractions in the background.
@@
  • We reinterpret visual model-based reinforcement learning from an information-theoretic perspective, pointing out the information conflict problem therein. Additionally, we discover the problem of positive-negative sample confusion when applying contrastive learning loss InfoNCE [29] in reinforcement learning.
@@
  • We evaluate MinCo on continuous control tasks with complex background distractions, and the results demonstrate that our method outperforms state-of-the-art model-based reinforcement learning methods.
@@ III-B Visual Model-Based Reinforcement Learning
The expectation is taken under the representation model and a dataset consisting of trajectories. Based on the learnd predictive dynamics model, Dreamer utilizes the actor-critic method to train policy by generating imagined transition data:
@@ IV-A Information Conflicts in Visual MBRL
$\geq\mathbb{E}[\ln\frac{p(s_{t}|o_{t},s_{t-1},a_{t-1})}{q(s_{t}|s_{t-1},a_{t-1})}]$
$-\text{KL}(p(s_{t}|s_{t-1},a_{t-1})\parallel q(s_{t}|s_{t-1},a_{t-1}))$
$=\mathbb{E}[\ln\frac{p(s_{t}|o_{t},s_{t-1},a_{t-1})}{p(s_{t}|s_{t-1},a_{t-1})}]$
$\equiv I(s_{t},o_{t}|s_{t-1},a_{t-1})$ (7)
@@ IV-B Negative-free Contrastive Learning for Visual MBRL
Essentially, these problems, particularly in the context of contrastive learning, arise from the improper selection of training samples and batches. Therefore, to address this issue, we need to construct appropriate training batches, and then adopt conflict-free training strategies. In this paper, we directly apply the negative-free contrastive learning algorithms, in which training batches only contain positive training samples, and hence avoid the problem of information conflicts and sample confusion. We adapt SimSiam [47] to visual MBRL as the objective of visual representation learning due to its simplicity and high learning efficiency.
@@ V-D How is the performance of MInCo compared to the state-of-the-art visual MBRL algorithms?
MInCo consistently achieves state-of-the-art performance on tasks in the distracted DeepMind Control suite benchmark.
@@
MInCo is robust against small subjects and dynamic backgrounds.
@@ Acknowledgments
Appendix A Proofs for Proposition 1 and 2

### v3

@@ Abstract
Model-based reinforcement learning, visual reinforcement learning, Continuous control under visual distractions, information conflicts.
@@ I Introduction
Among various forms of machine learning, reinforcement learning (RL), which models decision-making through trial and error, is one of the approaches most similar to human learning [1]. Deep reinforcement learning (DRL) utilizes deep neural networks to tackle high-dimensional tasks, achieving notable success in fields such as robotic manipulation [3, 4, 5], autonomous driving [6, 7, 8], video games [9], Go [10], and others [11]. Another fundamental human ability is planning actions based on predictions of the environment, which is especially useful in long-term tasks. Model-based reinforcement learning (MBRL), designed to mimic this ability, learns a dynamics model to predict the future and then performs policy optimization [12, 13, 14] or planning [15, 16, 17, 18] to determine the optimal action sequence. However, modeling environmental dynamics from high-dimensional visual observations remains challenging, particularly when faced with noise and distractions in the background.
@@
  • We reinterpret visual model-based reinforcement learning from an information-theoretic perspective, pointing out the information conflict problem therein.
@@
  • We evaluate MinCo on continuous control tasks with complex background distractions, and the results demonstrate that our method outperforms state-of-the-art model-based reinforcement learning methods. And we validate MInCo’s effectiveness in mitigating information conflicts through visualization.
@@ III-B Visual Model-Based Reinforcement Learning
The expectation is taken under the representation model and a dataset consisting of trajectories. Based on the learned predictive dynamics model, Dreamer utilizes the actor-critic method to train policy by generating imagined transition data:
@@ IV-A Information Conflicts in Visual MBRL
$\geq I(s_{t},o_{t}|s_{t-1},a_{t-1})$ (7)
@@ IV-B Negative-free Contrastive Learning for Visual MBRL
Essentially, these problems, particularly in the context of contrastive learning, arise from the improper selection of training samples and batches. Therefore, to address this issue, we need to construct appropriate training batches, and then adopt conflict-free training strategies. In this paper, we directly apply the negative-free contrastive learning algorithms, in which training batches only contain positive training samples, and hence avoid the problem of information conflicts and sample confusion. We adopt SimSiam [49] to visual MBRL as the objective of visual representation learning due to its simplicity and high learning efficiency.
@@ V-D How is the performance of MInCo compared to the state-of-the-art visual MBRL algorithms?
MInCo consistently achieves state-of-the-art performance on tasks in the distracted DeepMind Control suite benchmark
@@
MInCo is robust against small subjects and dynamic backgrounds
@@ Acknowledgments
Appendix A Proofs for Propositions 1 and 2